answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

When a single card is drawn from an ordinary 52 card deck find the probability of drawing a non-face card or a 7?

Since 7 is already a non-face-card, we know that there is 4 of each suit, and 3 royalty. 4(3) = 12, 52 - 12 = 40. 40/52 = 10/13.

If you roll a die 30 times how many times should you expect to roll a number less than 3?

The expected number of rolls under 3 is the number of rolls times the probability of rolling less than 3. So:

E(rolls less than 3 out of 30) = 30 * 1/3 = 30/3 = 10

How fast is emile heskey?

Heskey may be seen as a complete donkey. However this is a common mistake. Hekey is the fastest being in the whole world. Heskey's strength makes him the best player in the world. He keeps on scoring sweaty goals. Sweaty sweaty sweaty goals. he puts messi to shame. Heskey forever <3

When you Toss a coin 3 times and what is the probability of getting at least one head?

Pr(At least one head in 3 tosses) = 1 - Pr(No heads in 3 tosses)

= 1 - Pr(3 tails in three tosses) = 1 - [Pr(T)*Pr(T)*Pr(T)] since the three tosses are independent.

= 1 - 1/2 * 1/2 *1/2 = 1 - 1/8 = 7/8

What is the probability of 7 when tossing a set of fair dice one die at a time?

The 6 possible ways are:

1 and 6

2 and 5

3 and 4

4 and 3

5 and 2

1 and 6

6 combinations out of a possible 36 (6*6), therefore

6/36 or 1/6 or 1 in 6

What is the binomial probability rule?

  • Suppose you have n repetitions of an experiment with only two possible outcomes each time: success and failure,
  • Suppose the probability of success at any trial is independent of the outcomes of other trials and is a constant, p.

Then, the probability of exactly r successes (0 < r < n) is

[nCr]*p^r*(1-p)^(n-r) where nCr = n!/[r!*((n-r)!].

n! stands for n factorial [=fact(n) in Excel] and is equal to 1*2*...*n and 0! = 1.

Which approach to probability requires equally likely outcomes?

The modelling of a probability distribution function for an event from a theoretical approach.

What are the chances of a girl getting pregnant from putting it in underwater for not even 10 seconds?

Just as good as if she was not underwater.

If sperm come into proximity with the vagina, they will tend to migrate to the uterus and fallopian tubes and find the egg, if there is one. You don't even have to have penetration, and you don't even have to have ejaculation. Being underwater has nothing to do with it.

Is it true that probability predict the precise outcome of an individual event?

False. Probability only predicts the outcome. It does not assure the outcome.

What does the DIN standard mean to skiers?

it really means the binding of the skiing and the adjustment of it.

That makes no sense.

The DIN of a ski is very important to skiers because it is how easily the ski will release in a fall. The lower the number the easier the ski releases. Generally the lowest is around three and highest, pro racers, can be above 24. If the DIN is too low the ski can release when it is not supposed to, often if you hit a bump when going fast. An annoying crash at best, potential broken bones, or worse, at worst. If it is too high the skis will not release when they should. At best nothing happens, but this is very rare. Often you twist things, mostly knees but if you are unfortunate you can smash your skis into your face which really hurts or break bones by twisting your legs in ways they shouldn't.

Draw the probability tree diagram for flipping the coin two times showing the outcome?

Each toss outcome has a probability of 1/2; picture copied from the related link. The related link does a good job explaining tree diagrams and probabilities.

The odds against winning 5 in a lottery game are 179 to 2 What is the probability of winning?

If the odds are 179 to 2, the probability of winning is 0.01105, or; the chance of winning is 1.11%

A simple dice is tossed what is the probability of a two turning up?

There are 6 outcomes, a 2 is one of them so the probability is 1/6.